TOWARD A QUANTUM-MECHANICAL UNDERSTANDING OF THE HIGH-RESOLUTION INFRARED SPECTRUM OF CH+ 5

\underline{CHANDRA SAVAGE}, FENG DONG, DAVID J. NESBITT, JILA, National Institute of Standards and Technology and University of Colorado, Boulder, CO 80309.

Our recent measurement of the infrared spectrum of CH{+ \atop 5} in the C-H stretching region (2825-3100 cm^{−1}) has provided new insight into the behavior of this highly fluxional molecule. Examining four-line combination differences matched to 40 MHz, we have tentatively assigned two vibrational band origins near 2950 and 3025 cm^{−1}. Using the most complete ab initio calculations to date, these features can be assigned to symmetric and asymmetric stretches of the hydrogen-like CH_{2} component of CH{+ \atop 5}. Furthermore, we have assigned several ro-vibrational progressions within these bands. Fits of these lines to a simple spherical-top Hamiltonian yield a ground-state rotational constant of approximately 3.92 cm^{−1}, in agreement with theoretical predictions. With these initial results, the mystery of CH{+ \atop 5} is beginning to be unraveled.